Romnick is 3 more than twice the age of Levin. In 10 years, the sum of their ages will be 47. How old are they now?
Answer by macston(5194)

.
L=Levin's age; R=Romnick's age=2L+3yrs
.
(L+10yrs)+(R+10yrs)=47yrs
L+R+20yrs=47yrs
L+R=27yrs Substitute for R
L+(2L+3yrs)=27yrs
3L+3yrs=27yrs
3L=24yrs
L=8yrs Levin is currently 8 years old.
R=2L+3yrs=2(8yrs)+3yrs=16yrs+3yrs=19yrs Romnick is currently 19 years old.
.
CHECK:
(L+10yrs)+(R+10yrs)=47yrs
(8yrs+10yrs)+(19yrs+10yrs)=47yrs
18yrs+29yrs=47yrs
47yrs=47yrs